Classify each of these reactions as either acid-base, reduction-oxidation, or combustion.

a. ____CH3NH2 + ____O2 →____CO2 + ____H2O + ____N2

b. _____Zn(OH)2 + _____H3PO4 →_____Zn3(PO4)2 + _____H2O

c. _____WO3 + _____H2 →_____W + _____H2O

d. ____Al(OH)3 + ____HNO3 →____Al(NO3)3 + ____H2O

e. _____C5H12 + _____O2 →_____CO2 + _____H2O

f. ____AgBr + ____Cl2 → ____AgCl + ____Br2
Answer #1

lassify each of these reactions as either acid-base, reduction-oxidation, or combustion.

  1. ____CH3NH2 + ____O2 →____CO2 + ____H2O + ____N2

4 CH3NH2(g) + 9 O2(g) = 4 CO2(g) + 10 H2O(g) + 2 N2(g)

Reaction Type: Combustion

It is a high-temperature exothermic redox compound response between a fuel and an oxidant, normally air oxygen, that produces oxidized, regularly vaporous items, in a blend named as smoke. Burning in a fire delivers a fire, and the warmth created can make ignition self-maintaining. Ignition is frequently a muddled grouping of basic radical responses. Strong powers, for example, wood, first experience endothermic pyrolysis to create vaporous powers whose burning then supplies the warmth required to deliver a greater amount of them. Burning is frequently sufficiently hot that light as either gleaming or a fire is created. A basic illustration can be found in the burning of hydrogen and oxygen into water vapor, a response normally used to fuel rocket motors. This response discharges 242 kJ/mol of warmth and diminishes the enthalpy as needs be (at steady temperature and weight): Gour M is a burnable protest.

2H 2(g) + O 2(g) → 2H 2O(g)

Ignition of a natural fuel in air is constantly exothermic in light of the fact that the twofold bond in O2 is much weaker than other twofold bonds or combines of single bonds, and accordingly the arrangement of the more grounded bonds in the burning items CO2 and H2O brings about the arrival of energy. The bond energies in the fuel assume just a minor part, since they are like those in the burning items; e.g., the whole of the bond energies of CH4 is almost the same as that of CO2. The warmth of burning is around - 418 kJ per mole of O2 spent in the ignition response, and can be assessed from the essential piece of the fuel.[

===================================

  1. _____Zn(OH)2 + _____H3PO4 →_____Zn3(PO4)2 + _____H2O

3 Zn(OH)2 + 2 H3Po4 = Zn3(Po4)2 + 6 H2O
Reaction type: acid-base (double replacement)

It , otherwise called a twofold substitution response or metathesis, is a sort of synthetic response where two mixes respond, and the positive particles (cation) and the negative particles (anion) of the two reactants switch places, framing two new mixes or items

==========================

  1. _____WO3 + _____H2 →_____W + _____H2O

WO3 + 3 H2 = W + 3 H2O
Reaction Type: reduction-oxidation {Single Displacement (Substitution)]

A solitary uprooting response is a particular kind of oxidation-decrease response. A component or particle is supplanted by another in a compound. Single Displacement Reaction

==================================

  1. ____Al(OH)3 + ____HNO3 →____Al(NO3)3 + ____H2O

3 HNo3 + Al(OH)3 = Al(No3)3 + 3 H2O
Reaction type: acid-base (double replacement)

=======================================================================

  1. _____C5H12 + _____O2 →_____CO2 + _____H2O

5 C5H12 + 25 O2 = 25 CO2 + 3 H20
Reaction Type: reduction-oxidation {Single Displacement (Substitution)]

======================================================

f. ____AgBr + ____Cl2 → ____AgCl + ____Br2

2 AgBr + Cl2 = 2 AgCl + Br2
Reaction type: reduction-oxidation {Single Displacement (Substitution)]